# A phonon-driven mechanism for an emergent and reversible chirality in crystals

### v1

@@ Abstract
We demonstrate from first-principles calculations applied to the K3NiO2 crystal that a structural phase transition from an achiral to a chiral phase can be mediated by a soft zone boundary phonon and that it can be controlled by pressure and epitaxial strain. Calculations in the presence of an electric field hint at a rich physics dominated by the competition between chirality, polar, and achiral distortions. Induced by the interaction between chiral, polar, and axial phonons, an optimal parameter window for the permanent reversal of chirality is observed.
Periodic solids are chiral if their structures are described by space groups containing only proper symmetry operations that map, for example, a right-handed coordinate system onto a right-handed one, and not onto a left-handed one (“operations of the first kind”), [1] otherwise they are achiral [2]. In recent years there has been a growing interest in investigating certain chiral effects (both in real- and spin-space) in crystals such as the Chiral Induced Spin Selectivity (CISS) and spin spirals [3, 4], polar [5] and magnetic skyrmions [6, 7], topological states [8, 9] and chiral phonons [10, 11, 12] to mention a few. Noticeably, the crystallization process fixes the chirality and associated properties in most of these instances. While the concept of spontaneous symmetry breaking has so far been a fertile ground for important discoveries in condensed matter physics [13], with phenomena such as ferroelectricity, magnetism, or superconductivity, to date, it has not been applied to study the possibility of chirality spontaneously emerging from an achiral inorganic solid phase. Nevertheless, a chiral irreducible representation (IRREP) or order parameter can be formally defined [14], so that a group-subgroup relationship of the type achiral chiral between two phases is a physical possibility [15]. Despite the recent experimental reports about the spontaneous formation of gyrotropic domains observed in certain materials [16, 17, 18, 19, 20, 21, 22] below a critical temperature, many questions remain to be answered about their physical origin and properties. At the most fundamental level, we know little about which mechanisms stabilize a chiral order. Furthermore, we do not know if an enantiomeric excess mediated by external fields or other parameters in a nonpolar crystal and can be realized in the solid state.
In this work, we attempt to address these issues by exploring the achiral to chiral enantiomorphic phase transition observed in K3NiO2 (KNO) [23]. Using symmetry analysis and density functional theory (DFT) calculations (see the Supplementary file [24] for technical details and references [25, 26, 27, 28, 29, 30, 31, 32, 33] there included), we show that a soft zone boundary phonon mode triggers the emergence of chirality in this material with four well-like shapes in the energy landscape. We then explore how this structural chiral energy surface is enhanced or reduced by pressure or strain, and going further, we analyze how the chiral distortion can couple non-linearly with infrared (IR) and axial Raman active modes activated by an electric field. We prove that direct control over these polar and axial distortions is crucial to attain a permanent enantiomeric excess in a hysteresis-like manner and after said external field is removed. Our results can, therefore, be extrapolated to other chiral systems and allow us to explore strategies that may be adopted to optimize an enantiomeric excess in crystals via physical rather than possibly more invasive chemical procedures.
The high symmetry phase of KNO adopts an achiral tetragonal structure described by space group symmetry (no. 136). On cooling below 423 K, it undergoes a first-order phase transition to a low-temperature enantiomorphic phase of either 2 (no. 92) or 2 (no. 96) space groups [23]. Figure 1 shows the achiral high-temperature structure and illustrates the chiral distortion. Our symmetry analysis reveals that the 2/2 non-invariant distortion decomposes into symmetry adapted modes (SAMs) as , where Z4 and A1u are respectively the representation of the zone boundary () mode (that triggers the instability, as we will see below) and of an associated -point deformation which isotropy subgroup is 212. The mode involves the mostly in-plane motion of all the atom sites (although with larger K and O contributions in KNO) to form a spiral that preserves the 4-fold screw rotation of the parent phase. The pseudoscalar SAM describes a nonpolar K displacement along the direction. As a consequence of the O - K - O bond angle becomes a little smaller than 180∘ whilst the NiO units remain close to linear, consistent with experimental structure reports [23, 34, 35].
We perform our calculations with the collinear spins and the Hubbard-U correction (U = 4.2 eV) on Ni ions (required to open a band gap) [24]. We calculate the phonon spectrum in the phase and identify a doubly degenerate soft mode at the Z point with frequency = 1.35i THz and Z4 irreducible representation, consistently with the aforementioned SAM analysis. Despite the proposals in the past of different mechanisms for the emergence of chirality from an achiral reference - such as Jahn-Teller [36] as in CsCuClO3, or Peierls transitions [37] and orbital orders [38] as in the MgTi2O4 compound - in KNO the driving force thus appears to be solely of phononic nature. While other soft modes (at , X and M points [24]) are detected, the Z mode leads to the most stable low-symmetric configuration, as confirmed by the experiment. Since the Z4 soft boundary mode is doubly degenerate, we can build arbitrary linear combinations of eigenstates and calculate the energy as a function of their amplitude as illustrated in fig. 2. In particular, while the diagonal directions (a,a) and (a,-a) are associated with the realization of the P41212 and P43212 enantiomorphic domains, respectively, different settings lead to structures with different symmetries. In particular, the (a,0) and (0,b) directions produce degenerate orthorhombic phases with (no. 63) symmetry, which are achiral and optically inactive, while arbitrary (a,b) (with a0, b0 and ab) combinations generate systems with symmetry. The free energy F - which includes the chiral SAMs amplitudes - thus reads .
The electronic density of states (DOS) near the Fermi level (see Fig. 3(a)) is mainly populated by weakly hybridized O-/Ni- valence and K- conduction states with the ground state in a low-spin configuration. We find the density of states to be negligibly affected by the phase transition [24], which also agrees with the rather nominal calculated Born effective charges [24] and possibly suggests a short-range interaction mechanism for the emergence of chirality, in agreement with its zone boundary nature [39]. Concerning the behavior under mechanical perturbations, our results in Fig. 3(b)) show that even a tiny amount of hydrostatic pressure (above a value between 1.0 and 1.5 GPa) hardens the chiral mode and destroys the phase transition. This hardening under pressure may explain the observed decrease of the transition temperature upon replacing K with Rb or Cs, namely in Rb3NiO2 (Tc = 390 K) and Cs3NiO2 (achiral at room temperature) [34]. Conversely, an even more nuanced effect can be observed in the presence of strain. Indeed, the Z4 mode is respectively hardened or softened if the applied in-plane lattice deformation is tensile or compressive, as shown in Fig. 3(c). These results stem from a tension-compression asymmetry of the energy response in both the high and low symmetry phases. Nevertheless E() > E() always in the considered strain range (E is the energy, formula unit). Noticeably we find the instability, which is polar ( representation) and comes from short-range forces (due to the non-anomalous Born charges [24]), to persist under tensile strain, with E() = -2.3 meV upon = 3 %. While this value lies higher than the corresponding chiral state energy, it is possible that ferroelectricity may coexist with chirality at this or at a larger magnitude of . We leave this possibility open for future exploration.
Handedness flipping. Synthesized samples prepared by an azide-nitrate route resulted in racemic twinned crystals with approximately equal amounts of each enantiomorph (Flack parameter = 0.53) [23], illustrating the difficulty of preparing enantiomerically pure samples. At the same time, the free energy expansion in presence of both chiral of zone-centre pseudoscalar amplitudes gets modified as , which shows that the chiral order parameter can couple trilinearly to the A1u zone center mode. Therefore, any tensor containing the pseudoscalar irreducible representation can couple to the chiral Z4 mode as well through an invariant of the form and possibly induce control over the handedness of the system. One such tensor may be obtained by taking the product between vector () and axial () quantities. Indeed, the representation of this product can be decomposed into its irreducible components as (note that both the and products contain ), thus revealing a pseudoscalar component.
@@
We will now construct a model - based on the interaction between chiral, polar, and axial modes - that breaks the / degeneracy, realizes field-induced enantioselectivity and potentially creates a permanent enantiomeric excess. Applying an electric field results in the activation of polar in-plane and out-of-plane modes (with and representations respectively). The axial Raman active modes with representation can be switched on as well via trilinear interaction, following a scheme analogous to that of orthorhombic ErFeO3 [40]. In the following we will explore how these field-induced distortions couple with the Z4 instability. Along with the always present biquadratic couplings, the following terms are allowed:
@@
While is an enantioselective mechanism that requires the activation of the modes, breaks the (,0)/(0,) degeneracy of the achiral distortion. It is interesting to check if the energy barrier that separates the right- and left-handed domains can in practice be overcome by . This would lead to a permanent enantiomeric excess (i.e. surviving the 0 condition). To numerically test these ideas, we have extracted the relevant and modes via DFT structural relaxations in the presence of an electric field [41]. Importantly and since including collinear magnetism and the Hubbard-U interaction proves too numerically cumbersome in the presence of an electric field, we replace the K and Ni positions with Na and Au respectively to form a filled -shell paramagnetic Na3AuO2 (NAO) system with the same asymmetric unit of KNO and with a 2.1 eV DFT-computed band gap (without Hubbard correction). We note that although Na3AuO2 has been shown experimentally to adopt a slightly different structure [42], our "toy model" (see the SI [24] for the details of its construction) provides a helpful approximation of KNO. We apply an electric field E = 4.63x kV/m along the [1,1,1] direction to the high symmetry phase of Na3AuO2, perform a structural relaxation and extract the mixed polar-Raman distortion . Applied to the energy landscape, we observe that the overall effect of is to stabilize the (0,b) distortion while hardening all other orthogonal directions ((a,a) and (a,0)) as a consequence of the effect of [24]). Importantly, no energy difference between mirror-symmetric structures is detected. These features are also robust in the presence of an additional inducing compressive strain along the x-direction ( = 1 % [24]). The lack of right/left splitting can be explained if one realises that when [24]. Thus while the gyrotropic properties of KNO could be affected by the field, the absence of any or crossing prevents the realization of a permanent enantiomeric excess. Instead of exploiting the aforementioned nonlinear coupling in the presence of a constant bias along the [1,1,1] direction, we could use an axial external field - such as or [43, 44] (where defines an ionic displacement) - to switch the modes on. Assuming we can activate and phonons along the and directions respectively ( along the direction is set to zero), then 0 and the energy landscape may admit a sizable if the axial mode is large enough. For the sake of simplicity, we can assume to retain the polar-activated mode, although admitting a linear scaling of its amplitude to mimic an axial conjugate field. The effect of this strategy is shown in Fig. 4(a,b,c). Fixing the amplitude, a small polar mode as in Fig. 4(a) ( x 0.5) breaks the degeneracy between mirror equivalent enantiomers but cannot result in the crossing of the barrier. On the other hand, increasing the polar scaling amplitude to 1 shows in Fig. 4(b) that there exists an optimal window for the realization of a permanent enantiomeric excess, as measured by the difference between and concerning the distortion and reference. Finally, a larger amplitude (e.g. x 2) has the effect of stabilizing the polar state as reported in Fig. 4(c). Once the corresponding barrier has been crossed the system is expected to be remain scalemic (e.g. non racemic) in zero-field conditions. The overall handedness of the system can therefore be inverted, by flipping the sign of the electric field direction and with a magnitude consistent with fig. 4(b), consistently with a mechanism that is analogous to a hysteresis process.
Realizing a reversible and controllable enantiomeric excess is therefore interlaced with the direct activation of polar and axial modes. The absence of a true conjugate field for the chiral order indicates that any handedness-flipping interaction can activate achiral competing degrees of freedom as well (polar and in this case). While it is possible for a phonon state to be polar and chiral at the same time [21], so that the chirality and associated gyrotropic properties can be controlled by an electric field [45, 46] as in the case of the Pb5Ge3O11 compound [22], this requires the axial mode to be invariant [22] which is not the case for enantiomorphic crystals such as KNO. In practice realizing an enantiomeric excess (or "chirality poling") could be achieved via chiral phonon-light coupling. By symmetry, the tensor coupling with the Z soft eigenmode may take the form of the so-called "Lipkin zilch" [43], which defines the chirality density - - of a circularly polarised electromagnetic (EM) field. While no detailed study of an enantiomeric excess realized in the presence of a chiral phonon instability has been reported, the control over the phonon handedness in quartz has recently been obtained with the aid of circularly polarized X-rays[12].
Conclusions. In this work, we have explored the crystal chirality as a spontaneous order driven by a phonon instability with the K3NiO2 compound as a model. In particular, a high-symmetry phase is lowered by a degenerate point soft mode to a pair of enantiomorphic structures with and symmetries. We have shown that an epitaxial strain can be used to either soften or harden the unstable mode, while a moderate hydrostatic pressure stabilizes the achiral phase. Finally, we have explored the possible realization of an enantiomeric excess in KNO mediated by a coupling between chiral, polar, and axial orders. Our findings highlight the rich phenomenology triggered by a static electric field, including a competition between chirality and orthorhombic () phases and enantioselectivity. Nevertheless, we have numerically shown that the modes, when activated via polar-polar-Raman coupling [40], cannot efficiently lift the right/left degeneracy, so realizing a permanent handedness flipping could more likely be achieved by using an axial conjugate field (on top of an electric field) such as rotating EM fields or structural deformations. Additional theoretical and experimental analysis would be appealing to further understand our predictions on phonon-mediated chirality and the related new perspectives for enantioselectivity.

### v2

@@ Abstract
We demonstrate through first-principles calculations applied to the K3NiO2 crystal that a structural phase transition from an achiral to a chiral phase can be mediated by a degenerate soft phonon mode and controlled by pressure and epitaxial strain.
Breaking such degeneracy with an electric field generates a competition between enantiomorphic, polar, and orthorhombic displacements.
Originated by the interaction between spontaneous chiral and induced polar and axial modes, an optimal parameter window for converting the handedness of the system into its opposite kind
is observed.
Periodic solids are chiral if their structures are described by space groups containing only proper symmetry operations that map, for example, a right-handed coordinate system onto a right-handed one, and not onto a left-handed one (“operations of the first kind”), [1] otherwise they are achiral [2].
In recent years there has been a growing interest in investigating certain chiral effects (both in real- and spin-space) in crystals such as the Chiral Induced Spin Selectivity (CISS) and spin spirals [3, 4], polar [5] and magnetic skyrmions [6, 7], topological states [8, 9] and chiral phonons [10, 11, 12] to mention a few.
Noticeably, the crystallization process fixes the chirality and associated properties in most of these instances.
While the concept of spontaneous symmetry breaking has so far been a fertile ground for important discoveries in condensed matter physics [13], with phenomena such as ferroelectricity, magnetism, or superconductivity, to date, it has not been applied to study the possibility of chirality spontaneously emerging from an achiral solid phase.
Nevertheless, a chiral irreducible representation (IRREP) or order parameter can be formally defined [14], so that a group-subgroup relationship of the type achiral chiral between two phases is a physical possibility [15].
Despite the recent experimental reports about the spontaneous formation of gyrotropic domains observed in certain materials [16, 17, 18, 19, 20, 21, 22] below a critical temperature,
many questions remain to be answered about their physical origin, properties and the way the can be controlled. Importantly, it is not clear if the handedness of a chiral material can be effectively converted into the opposite kind via the sole application of external fields.
In this work, we attempt to address these issues
by exploring the achiral to chiral enantiomorphic phase transition observed in K3NiO2 (KNO) [23].
Using symmetry analysis and density functional theory (DFT) calculations (see the Supplementary file [24] for technical details and references [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] there included), we show that a soft zone boundary phonon mode triggers the emergence of chirality in this material with four well-like shapes in the energy landscape.
We then explore how this structural chiral energy surface is enhanced or reduced by pressure or strain, and going further, we analyze how the chiral distortion can couple non-linearly with infrared (IR) and axial Raman active modes activated by an electric field.
We prove that direct control over these polar and axial distortions is crucial to attain a permanent enantiomeric excess in a hysteresis-like manner and after said external field is removed.
We discuss how this mechanism can be generally applied to several families of materials, with having one degenerate instability generating both opposite enantiomers as the sole requirement.
Figure 1: Scheme of the soft chiral phonon distortion in K3NiO2. The purple, green, and red spheres correspond to K, Ni, and O atoms. The ions are in the configuration (1x1x2 supercell), while the golden arrows and the red spirals represent the chiral ( left , right) displacement (for clarity, only the main K-motion is shown).
The high symmetry phase of KNO adopts an achiral tetragonal structure described by space group symmetry (no. 136).
On cooling below 423 K, it undergoes a first-order phase transition to a low-temperature enantiomorphic phase of either 2 (no. 92) or 2 (no. 96) symmetry [23]. Our analysis of the experimentally reported structures [23] hence shows that the corresponding symmetry-lowering structural deformation decomposes into a main zone boundary (Z = 0,0,1/2) and a secondary pseudoscalar zone centre symmetry adapted modes (SAMs) with Z4 and A1u representation, respectively.
We perform our calculations with the collinear spins and the Hubbard-U correction (U = 4.2 eV, computed via the Cococcioni’s method [28]) on Ni ions (required to open a band gap) [24] and compute [34, 35] the phonon spectrum in the phase. We identify, consistently with the previous SAM analysis, a doubly degenerate soft mode at the Z point with frequency = 1.35i THz and Z4 irreducible representation.
While other soft modes (at , X and M points [24]) are detected, the Z mode leads to the most stable low-symmetric configuration, as confirmed by the experiment [36, 23, 37]. The computed distortion alone induces the chiral phases and chiefly involves the in-plane displacement (0.4 Å/ion) of potassium ions from their position at the Ni-O-K-O linear chains to form a handed 4{1,3} spiral as in fig. 1. As a consequence, the chain O - K - O bond angle becomes smaller than 180∘ whilst the NiO units remain close to linear, in agreement with experimental structure reports [23, 36, 38].
A secondary A1u SAM, observed upon optimization of the chiral structures, is instead solely constituted by a small displacement (0.006 Å/ion) of tetrahedrally coordinated K atoms along the axis (see supplementary section III [24]).
Since the Z4 soft boundary mode is doubly
degenerate, we can build arbitrary linear combinations of eigenstates and calculate the energy as a function of their amplitude as illustrated in fig. 2.
The free energy F - which includes the chiral SAMs amplitudes - thus reads .
In particular, while the diagonal directions (,) and (, -) are associated with the realization of the 212 and 212 enantiomorphic domains, respectively, different settings lead to structures with different symmetries. More specifically, the (,0) and (0,) directions produce degenerate orthorhombic phases with (no. 63) symmetry, which are achiral and optically inactive, while arbitrary (,) (with 0, 0 and ) combinations generate systems with symmetry.
Figure 2: Schematic 2D projection of the surface energy given by condensing the two-dimensional mode (amplitudes in arbitrary units).
The central black circle corresponds to the high symmetry phase.
The (no. 63) energy minimum points
are highlighted with white circles.
Red and pink-filled circles indicate the degenerate ground state points in the (no. 96) and (no. 92) enantiomorphic phases, respectively.
The electronic density of states (DOS) near the Fermi level (see Fig. 3(a)) is mainly populated by weakly hybridized O-/Ni- valence and K- conduction states with the ground state in a low-spin configuration.
We find the density of states to be negligibly affected by the phase transition [24], which also agrees with the rather nominal calculated Born effective charges [24] and possibly suggests a short-range interaction mechanism for the emergence of chirality, in agreement with its zone boundary nature [39]. Concerning the behavior under mechanical perturbations, our results in Fig. 3(b)) show that even a tiny amount of hydrostatic pressure (above a value between 1.0 and 1.5 GPa) hardens the chiral mode and destroys the phase transition.
This hardening under pressure may explain the observed decrease of the transition temperature upon replacing K+ with larger Rb+ or Cs+ ions, namely in Rb3NiO2 (Tc = 390 K) and Cs3NiO2 (achiral at room temperature) [36].
Conversely, an even more nuanced effect can be observed in the presence of strain.
Indeed, the Z4 mode is respectively hardened or softened if the applied in-plane lattice deformation is tensile or compressive, as shown in Fig. 3(c).
These results stem from a tension-compression asymmetry of the energy response
in both the high and low symmetry phases.
Nevertheless E() > E() always in the considered strain range (E is the energy, formula unit).
Noticeably we find the instability, which is polar ( representation) and comes from short-range forces (due to the non-anomalous Born charges [24]), to persist under tensile strain, with E() = -2.3 meV upon = 3 %.
While this value lies higher than the corresponding chiral state energy, it is possible that ferroelectricity may coexist with chirality at this or at a larger magnitude of .
We leave this possibility open for future exploration.
Figure 3: (a) Partial electronic density of states near the Fermi level (P42/mnm phase). The dashed green curve represents K states and the grey curve the total DOS. (b) and (c) E() behavior as a function of the hydrostatic pressure and epitaxial in-plane strain respectively.
Handedness conversion.
Synthesized samples prepared by an azide-nitrate route resulted in racemic twinned crystals with approximately equal amounts of each enantiomorph (Flack parameter = 0.53) [23], illustrating the difficulty of preparing enantiomerically pure samples.
At the same time, the Z4 and A1u representations are coupled by the invariant.
Therefore, any tensor containing the pseudoscalar IRREP interacts with the chiral Z4 mode through an invariant of the form , possibly allowing control over the handedness of the system.
One such tensor may be obtained by taking the product between vector () and axial () quantities [40].
Indeed, the representation of this product can be decomposed into its irreducible components as (note that both the and products contain ), thus revealing a pseudoscalar component.
@@
We will now construct a model - based on the interaction between chiral, polar, and axial modes - that breaks the / degeneracy, realizes field-induced enantioselectivity and potentially creates a permanent enantiomeric excess.
Applying an electric field results in the activation of polar in-plane and out-of-plane modes (with and representations respectively).
The axial Raman active modes with representation can be switched on as well via the trilinear interaction, following a scheme analogous to that of orthorhombic ErFeO3 [41].
In the following we will explore how these field-induced distortions couple with the Z4 instability.
Along with the always present biquadratic couplings, the following terms are allowed:
@@
Figure 4: DFT-evaluated energy landscape as a function of the amplitude in Na3AuO2 and in the presence of additional in-plane polar and axial independent distortions (continuous segments are a guide to the eye).
In this setting, = 0.
The magnitude is fixed to be five times the DFT-obtained value (via -- coupling, see main text),
while the amplitude is scaled (0.5, 1 and 2 in figs. (b), (c) and (d) respectively). Fig.(b) shows the onset of a handedness flipping via crossing the energy barrier.
While is an enantioselective mechanism that requires the activation of the modes, breaks the
(,0)/(0,) degeneracy of the achiral distortion.
It is interesting to check if the energy barrier that separates the right- and left-handed domains can in practice be overcome by .
This would lead to a permanent enantiomeric excess (i.e. surviving the 0 condition).
To numerically test these ideas, we have extracted the relevant and modes via DFT structural relaxations in the presence of an electric field [42].
Importantly and since including collinear magnetism and the Hubbard-U interaction proves too numerically cumbersome in the presence of an electric field, we replace the K and Ni positions with Na and Au respectively to form a filled -shell paramagnetic Na3AuO2 (NAO) system with the same asymmetric unit of KNO and with a 2.1 eV DFT-computed band gap (without Hubbard correction).
We note that although Na3AuO2 has been shown experimentally to adopt a slightly different structure
[43], our "toy model" (see the SI [24] for the details of its construction) provides a helpful approximation of KNO.
We apply an electric field E = 4.63x kV/m along the [1,1,1] direction of the high symmetry phase of Na3AuO2, perform a structural relaxation and extract the mixed polar-Raman distortion .
Applied to the energy landscape, we observe that the overall effect of is to stabilize the distortion while hardening all other orthogonal directions ((a,a) and (a,0)) as a consequence of the effect of [24]).
Importantly, no energy difference between
mirror-symmetric structures is detected.
These features are also robust in the presence of an additional inducing compressive strain along the -direction ( = 1 % [24]).
The lack of right/left splitting can be explained if one realises that when [24].
Thus while the gyrotropic properties of KNO could be affected by the field, the absence of any or crossing prevents the realization of a permanent enantiomeric excess. Instead of exploiting the aforementioned nonlinear coupling in the presence of a constant bias along the [1,1,1] direction, we could use an axial external field - such as or [44, 40] (where defines an ionic displacement) - to switch the modes on.
Assuming we can activate and phonons along the and directions respectively ( along the direction is set to zero), then 0 and the energy landscape may admit a sizable if the axial mode is large enough.
For sake of simplicity, we can assume to retain the polar-activated mode, although admitting a linear scaling of its amplitude to mimic an axial conjugate field.
The effect of this strategy is shown in Fig. 4(a,b,c).
Fixing the amplitude, a small polar mode
as in Fig. 4(a)
( x 0.5) breaks the degeneracy between mirror equivalent enantiomers but cannot result in the crossing of the barrier.
On the other hand, increasing the polar scaling amplitude to 1 shows in Fig. 4(b) that there exists an optimal window for the realization of a permanent enantiomeric excess, as measured by the difference between and concerning the distortion and reference.
Finally, a larger amplitude (e.g. x 2) has the effect of stabilizing the polar state as reported in Fig. 4(c).
Once the corresponding barrier has been crossed the system is expected to be remain scalemic (e.g. non racemic) in zero-field conditions.
The overall handedness of the system can therefore be inverted, by changing the sign of the electric field direction and with a magnitude consistent with fig. 4(b), consistently with a mechanism that is analogous to a hysteresis process.
Discussion. The proposed model can be applied to several other classes of materials as long as chiral orders with opposite handedness emerge from one multidimensional IRREP, a condition present in many other known systems. For instance, a (0,0,1/3) SAM induces a transition in hexagonal halide perovskites such as CsCuCl3 [45, 46]. At the same time, spinels such as MgTi2O4 [47, 48, 49, 50] along with SiO2 cristobalite [51, 52, 53, 54] show a displacive transition triggered by a soft X4 (0,1,0) mode.
These and X4 IRREPS span both enantiomorphic and orthorhombic displacements and calling
(,) the order parameter ( = , X4), we can generally define V and V with the same features of eqs. 1 and 2 (see SI [24], section V).
Additionally, several X4 spinels may adopt cation-ordered structures of symmetry [55] (e.g. LiZnNbO4 [56, 57, 58], ratio = 1:1), a process suitable to be influenced by V and V due to the X3 character of the occupational order parameter.
Realizing a reversible and controllable enantiomeric excess is therefore interlaced with the direct activation of polar and axial modes.
The absence of a true conjugate field for the chiral order indicates that any handedness-flipping interaction can activate achiral orthorhombic competing degrees of freedom as well.
While it is possible for a symmetry mode to be polar and chiral at the same time [21], so that the chirality and associated gyrotropic properties can be controlled by an electric field [59, 60] as in the case of Pb5Ge3O11 [22],
this requires the axial mode to be invariant [22] which is not the case for enantiomorphic crystals such as KNO.
In practice realizing an enantiomeric excess (or "chirality poling") could be achieved via chiral phonon-light coupling and, possibly, by working close to the transition temperature via adiabatic cooling.
By symmetry, the tensor coupling with the Z soft eigenmode may take the form of the so-called "Lipkin zilch" [44], which defines the chirality density - - of a circularly polarised electromagnetic (EM) field.
While no detailed study of an enantiomeric excess realized in the presence of a chiral phonon instability has been reported, the control over the phonon handedness in quartz has recently been obtained with the aid of circularly polarized X-rays[12].
Conclusions.
We have observed ab initio how the chirality in crystals can emerge as a displacive spontaneous order triggered by a degenerate soft mode. By using the K3NiO2 compound as a model system, we have shown how its symmetry is lowered from to the enantiomorphic , pair by a soft Z4 mode which can be controlled by epitaxial strain and hydrostatic pressure.
Finally, we have numerically explored the possible realization of a stable enantiomeric excess mediated by a coupling between chiral, polar, and axial orders.
Our model applies as long as order parameters with opposite handedness belong to the same representation, a condition encountered in a large variety of systems.
Importantly, we have shown that the modes, when activated via polar-polar-Raman coupling [41], cannot efficiently lift the right/left degeneracy and favor an orthorhombic distortion ( point group) instead.
Hence, realizing a permanent handedness conversion could more likely be achieved by using an axial conjugate field on top of an electric field.
Additional analysis would be appealing to understand how the gyrotropic physical properties of crystals are affected by this perspective on chirality.